A rectangular field is to be enclosed and divided into four equal lots by fences parallel to one of the sides. a total of 10,000 meters of fence are available. Find the area of the largest field that can be enclosed.


1
Expert's answer
2022-05-16T17:53:07-0400

The total length will be x


and the height will be y


Needed Equations:


Perimeter of this diagram


10000=2x+5y


Total Area A=xy


Solve for y using the equation


10000=2x+5y


5y=10000-2x


y=2000-2/5x


Substitute the equation for y into the function for area.


A=x(2000-2/5x)=2000x-2/5x2

Find the derivative of the equation for area.


A'=2000-4/5x


Use the derivative equation in order to find the critical point(s) that maximize the area.


Critical points are when


A'=0 and when A' does not exist. It is also good to check the endpoints of an equation in order to check for a maximum or minimum.Since A' always exists, only find where A'=0


(there will be no endpoints to check since this is a field).


0=2000 -4/5x


x=2000/(4/5)=2500


A'is positive when x<2500 and A' is negative when x>2500 , therefore meaning that x=2500 is a maximum. Since this value is a maximum, the area is maximized when the total length is 2500 m.


Find the height ( y ) when x=2500


y=2000-2/5x=2000-2/5(2500)=1000 m


The dimensions that will maximize the area the total area of the pig pen will be


2500m by 1000 m


A=2500x1000=2.5x105 m2